S12-6 (similar to) Question Help Your grandmother would like to share some of her fortune with you. She offers to give you moi Present Value of Annuity of $1 Present Value of Annuity of $1 Periods 4% 10% 12% 14% 16% 18% 1% 2% 3% 5% 6% 8% 20% 0.862 1.X i Future Value of Annuity of $1 - Future Value of Annuity of $1 Periods 1% 2% 3% 4% 5% 6% 8% 10% 12% 14% 16% 18% 20% 1.000- X i Present Value of $1 Present Value of $1 5% 14% 18% 20% Periods 1% 2% 3% 4% 6% 8% 10% 12% 16% 0.877 0.769 0.990 0.962 0.iFuture Value of $1 Future Value of $1 2% 3% 4% 1.020 5% 6% 1030 1040 1050 1 060 1.080 1.100 1.120 8% 10% 12% 14% 16% 18%| 20
Answer #1
scenario 1
1) 8750 a year at the end of each of the next eight year
present value = 8750*5.747 = 50286.25
2) lumpsum now = 48750
3) 99350 (lump sum) eight year from now
present value = 99350*0.540 = 53649
Highest net present value = 53649
scenario 2
1) 8750 a year at the end of each of the next eight year
present value = 8750*5.335 = 46681.25
2) lumpsum now = 48750
3) 99350 (lump sum) eight year from now
present value = 99350*0.467 = 46396.45
Highest present value = 48750